# SACNet: A Spatially Adaptive Convolution Network for 2D Multi-organ Medical Segmentation thanks: *Corresponding author.This work is supported by National Natural Science Foundation of China [No. 62371156] and Natural Science Foundation of Guangdong Province [No. 2022A1515011629]

### v1

@@ Abstract
Multi-organ segmentation in medical image analysis is crucial for diagnosis and treatment planning. However, many factors complicate the task, including variability in different target categories and interference from complex backgrounds. In this paper, we utilize the knowledge of Deformable Convolution V3 (DCNv3) and multi-object segmentation to optimize our Spatially Adaptive Convolution Network (SACNet) in three aspects: feature extraction, model architecture, and loss constraint, simultaneously enhancing the perception of different segmentation targets. Firstly, we propose the Adaptive Receptive Field Module (ARFM), which combines DCNv3 with a series of customized block-level and architecture-level designs similar to transformers. This module can capture the unique features of different organs by adaptively adjusting the receptive field according to various targets. Secondly, we utilize ARFM as building blocks to construct the encoder-decoder of SACNet and partially share parameters between the encoder and decoder, making the network wider rather than deeper. This design achieves a shared lightweight decoder and a more parameter-efficient and effective framework. Lastly, we propose a novel continuity dynamic adjustment loss function, based on t-vMF dice loss and cross-entropy loss, to better balance easy and complex classes in segmentation. Experiments on 3D slice datasets from ACDC and Synapse demonstrate that SACNet delivers superior segmentation performance in multi-organ segmentation tasks compared to several existing methods.
@@ I Introduction
Multi-organ medical image segmentation, which simultaneously provides clearer visualization of anatomical and pathological structures across multiple organs, greatly enhances the efficiency and accuracy of diagnosis [1]-[2].
@@
However, it remains challenging due to the following difficulties: (1) Variability in different target categories. As shown in Fig. 1, not only do different organs in the same slice dataset vary significantly in size, shape, and texture, but different slices of the same organ also differ greatly. This variability makes it difficult for segmentation models to generalize effectively, as they may overfit to specific features seen during training and struggle with unseen morphological variations. (2) Interference from complex backgrounds. Fig. 1 shows the complexity of the surrounding anatomical structures and the presence of varying textures, intensities, and overlapping tissues, which introduce substantial difficulties. These intricate backgrounds can lead to ambiguous boundaries and obscure the precise contours of the target organs. Consequently, the model’s ability to precisely identify and segment the organs is compromised, often resulting in incomplete or inaccurate segmentation outcomes.
Figure 1: Overview of challenges in multi-organ segmentation tasks and the effect of different convolutional techniques. The above figure: challenges faced in multi-organ segmentation tasks. The four images below: effects of standard convolutional kernel, dilated convolutional kernel, large-kernel convolutional kernel, and standard deformable convolutional kernel.
Recently, most existing studies [3]-[16] have utilized standard convolutional kernels, dilated convolutional kernels, large-kernel convolutional kernels, transformers, or hybrid methods as building blocks to design specific network architectures to improve segmentation accuracy. As shown in Fig. 1, dilated and large-kernel convolutional kernels have a larger receptive field compared to standard convolutional kernels. The transformers [9] capture long-range dependencies more effectively through self-attention mechanisms. However, due to significant differences between different organs, these models lack an inherent mechanism to adjust the specific receptive field according to different segmentation targets, inevitably losing the perception of the corresponding structures. In this context, some researchers [17]-[20] have started to introduce deformable convolutions and deformable attention transformers for multi-organ medical image segmentation. However, deformable convolutions [17] have limited flexibility in sampling points and high computational complexity. Deformable attention transformers [19] perform poorly when handling imbalanced label data, and the need for large amounts of data for generalization poses significant challenges for medical image tasks, where data is relatively scarce. Some researchers have also attempted to improve loss functions to better adapt to multi-organ medical image segmentation tasks. The standard Dice loss, due to cosine similarity’s disregard for data distribution characteristics [26], treats each class equally, which tends to result in imbalanced segmentation. Various loss functions [21]-[26] based on dice loss have been proposed to further reinforce the stringent constraints on segmentation. Building on this foundation, structured losses combined with continuous dynamic adjustment mechanisms will potentially further improve the accuracy of segmentation.
To tackle the above obstacles, we propose a novel framework, SACNet, involving an Adaptive Receptive Field Module (ARFM), widenet strategy, and dynamic continuity adjustment loss function. (1) To address the variability in different target categories and the interference from complex backgrounds, we propose the ARFM, which combines DCNv3 with a series of tailored block-level and architecture-level designs similar to transformers to enhance the perception of the geometric structure by adaptively focusing on different segmentation targets. Consider that DCNv3 allows the network to freely learn geometrical changes, which can cause perceptual regions to wander; therefore, our ARFM adopts the strategy of grouped convolutions. It divides the bias generation process of DCNv3 into groups corresponding to the segmentation target categories, with each group adaptively adjusting the receptive field according to different targets. Additionally, ARFM introduces the Feed-Forward block to capture specific patterns in the different segmentation targets. It also introduces the Layer Scale block and the DropPath block to prevent the model from excessively freely learning geometric changes. (2) To address the challenge of high memory and computation complexity, we adopt the WideNet strategy [31]. We share the DCNv3 block projection weights of ARFM modules between the encoder and decoder when generating the original location-aware modulation scalars and offsets. Through the FFN block, SACNet can train more parameters in width; by sharing parameters between DCNv3 blocks, SACNet can train fewer parameters in depth. In this method, the network goes wider instead of deeper, achieving a shared lightweight decoder and a more parameter-efficient and effective framework. (3) To address the problem of imbalanced segmentation caused by imbalanced label data, we propose a novel Continuity Dynamic Adjustment Loss Function (CTLoss). Based on t-vMF Dice loss and cross-entropy loss, CTLoss adaptively adjusts using the Intersection Over Union (IOU) of each segmentation category on the validation set after each epoch. Our CTLoss automatically adjusts to use more compact similarities for easy classes and wider similarities for difficult classes during training.
To summarize, our contributions are as follows:
  • We propose an Adaptive Receptive Field Module to enhance the perception of geometric structures by adaptively focusing on different segmentation targets.
  • We design a novel convolutional network that expands the network’s width rather than its depth, achieving a shared lightweight decoder and a more parameter-efficient and effective framework.
  • We propose a novel continuity dynamic adjustment loss function based on IOU, which better balances easy classes and difficult classes.
  • Our proposed SACNet method achieves state-of-the-art performance on two different multi-organ medical image segmentation datasets, as demonstrated by various ablation studies highlighting its effectiveness.
II RELATED WORK
II-A Methods based on Network Design
Various methods [3]-[16] have been proposed to achieve better performance by designing specific network structures and modules for multi-organ segmentation. (1) Methods based on the convolutional kernel and transformers: Since the introduction of UNet [3], Convolutional Neural Networks (CNNs) have become the most common choice. To enhance the limited receptive field, researchers [5] have attempted to replace standard convolutions with dilated convolutions and large-kernel convolutions. With the success of vision transformers, TransUNET [9] and similar methods [10]-[12] integrate transformers with convolutions as building blocks for the encoder. SwinUNet [13] and similar methods [14]-[16] aim to create a fully transformer-based model. (2) Methods based on deformable convolution and deformable transformers: Recently, deformable convolutions have been applied to medical image segmentation [17]-[18]. [27] extended standard deformable convolutions to DCNv2, and DCNv3 [28] further enhanced feature extraction at different scales by building on DCNv2. AgileFormer [20] introduced the deformable attention transformer, integrating deformable convolutions into patch embedding, self-attention, and positional encoding within the model. Considering the limitations of convergence difficulty and perceptual region wander caused by excessive randomness, our work integrates the knowledge of deformable convolution v3 and designs the ARFM as a building block for the encoder. The ARFM can adaptively adjust the receptive field according to different segmentation targets and mitigate the excessive flexibility of DCNv3, stably enhancing the perception of different segmentation regions in the feature extraction process.
Figure 2: Overview of the SACNet architecture. The left side of the network represents the encoder, which uses four stages of adaptive spatial aggregation feature extraction, while the right side shows the shared lightweight decoder that generates the segmentation maps. Each block of the ARFM employs pre-trained DCNv3 that adaptively aggregates spatial features. The Feed-Forward block captures specific patterns in the different segmentation targets and introduces the Layer Scale block and the DropPath block to prevent the model from excessively freely learning geometric changes.
II-B Methods based on Loss Function
In recent studies, various loss functions based on Dice loss have been proposed to further improve the Dice Similarity Coefficient (DSC) [21]-[25]. Li et al. [21] proposed a generalized Dice loss that utilizes the class re-balancing properties of the generalized Dice overlap, achieving a robust and accurate deep learning loss function for unbalanced segmentation. Shit et al. [22] introduced a similarity measurement termed centerline Dice, which is calculated based on the intersection of the segmentation masks and the skeleton. Wang et al. [23] improved Dice loss by mining the information of negative areas. Kato and Hotta [25] applied the t-vMF similarity to the Dice loss, adjusting the compactness of the similarity for each class. Drawing inspiration from this, our work proposes a novel Continuity Dynamic Adjustment Loss Function (CTLoss), which combines t-vMF dice loss and cross-entropy loss. Our CTLoss introduces a new dynamic updating mechanism, updating with the average IOU of each class on the validation set after each epoch, better guiding the network to balance easy classes and difficult classes.
III METHOD
III-A Encoder
Recent advancements in large-scale models [27] based on Convolutional Neural Networks (CNNs) have shown the potential of Deformable Convolution v3 (DCNv3), which has demonstrated greater effectiveness in various downstream visual tasks. In our SACNet model, we primarily utilize the ARFM module, which employs the pre-trained DCNv3 block as its core operation. This module integrates the DCNv3 block with a series of tailored block-level and architecture-level designs similar to Transformers to serve as the building blocks of the spatially adaptive pyramid feature encoder. Specifically, our input is an image size of 3×224×224, starting with a Stem Layer, which applies convolution followed by layer normalization and activation functions. Next, the input passes through four stages with embedding sizes of 112, 224, 448, and 896, respectively. Each stage consists of a different number of ARFM modules and a down-sampling module, with each ARFM module composed of five separate blocks: the DCNv3 block, which provides the model with the capability to adjust specific adaptive receptive fields according to different segmentation targets; the Layer Norm block, which performs normalization on the input samples; the Feed-Forward block, which captures specific patterns in the different segmentation targets; the Layer Scale block and DropPath block, which prevent the model from excessively freely learning geometric changes. The implementation of ARFM is shown in Equations (1)-(3) as follows:
$DCN(p_{0})=\sum_{g=1}^{G}\sum_{k=1}^{K}w_{g}m_{gk}x_{g}(p_{0}+p_{k}+\Delta p_{gk}),$ (1)
$z_{l}^{\prime}=DP\left(\gamma_{1}\cdot N\left(DCN\left(z_{l-1}\right)\right)+z_{l-1}\right),$ (2)
$z_{l}=DP\left(\gamma_{2}\cdot N\left(FFN\left(z_{l}^{\prime}\right)\right)+z_{l}^{\prime}\right),$ (3)
where $K$ represents the total number of sampling points, and $k$ enumerates each sampling point. $G$ denotes the total number of aggregation groups. For the $g$-th group, $\mathbf{w}_{g}\in\mathbb{R}^{C\times C^{\prime}}$ represents the location-irrelevant projection weights, where $C^{\prime}=\frac{C}{G}$ defines the group dimension. $m_{gk}\in\mathbb{R}$ denotes the modulation scalar of the $k$-th sampling point in the $g$-th group, normalized by the softmax function along dimension $K$. $\mathbf{x}_{g}\in\mathbb{R}^{C^{\prime}\times H\times W}$ represents the sliced input feature map. $p_{k}$ denotes the $k$-th location of the pre-defined grid sampling $\{(-1,-1),(-1,0),\ldots,(0,+1),\ldots,(+1,+1)\}$ as in regular convolutions. $\Delta\mathbf{p}_{gk}$ is the offset corresponding to the grid sampling location $p_{k}$ in the $g$-th group. $DCN$ stands for deformable convolution v3. $\gamma_{1}$ and $\gamma_{2}$ are the inter-layer scaling factors, $N$ denote layer normalization, and $DP$ denotes the droppath function.
The third layer of the encoder typically captures better feature representations than the others [28]. Through extensive experiments, we ultimately distribute more computation to the third encoder stage with an ARFM module ratio of [4, 4, 21, 4], respectively. A skip connection is utilized inside each stage to address the issue of vanishing gradients and facilitate information flow. Our encoder generates four feature representations of size $\{1\times(112\times 2^{i})\times\frac{H}{2^{2+i}}\times\frac{W}{2^{2+i}}\}$ in four different stages where $i=\{0,1,2,3\}$, which are passed to the decoder through skip connections. All modules are connected in the sequence shown in Fig. 2.
III-B Decoder
The SACNet decoder has four consecutive decoder blocks and one bottleneck block. Each decoder block starts with layer normalization, and then upsamples the inputs by a factor of two to match the size of the encoder outputs. Next, we apply the ARFM module to reshape the features, followed by layer normalization while keeping the input dimension the same. We then concatenate it with the encoder output through the skip connection. Similar to regular convolution, when the DCNv3 block generates offsets $\Delta{p}_{gk}$ and modulation scalars $m_{gk}$, different convolutional neurons in different groups have independent linear projection weights. Therefore, as the total number of sampling points increases, the model’s parameter and memory complexity grow rapidly, significantly limiting the model’s efficiency. To address this problem, we borrow the idea from WideNet [31], which shares some model parameters in depth. We share the DCNv3 block projection weights of ARFM modules between the encoder and decoder when generating the original location-aware modulation scalars and offsets. Through the Feed-Forward blocks, SACNet can train more parameters in width; by sharing parameters between DCNv3 blocks, SACNet can train fewer parameters in depth. Then, we apply bilinear interpolation to resize the images to $H\times W\times N$, where $N$ is the number of classes, used as the final segmented maps of the decoder, and pass through Softmax to classify each pixel. The operation of the bottleneck layer is the same as the other four decoder blocks, except that the bottleneck gathers more information in the channel of the image while the image size remains the same. Additionally, our decoder does not incorporate any extra techniques or strategies such as residual connections [29] or deep supervision [30].
Figure 3: Qualitative comparisons of the segmentation performance of our approach alongside other segmentation methods. The first and second rows show the results of the Synapse dataset, where the color codes are: yellow for the aorta, green for the gallbladder, orange for the left kidney, cyan for the right kidney, plum for the liver, purple for the pancreas, indigo for the spleen, and crimson for the stomach. The third and fourth rows show the results of the ACDC dataset, where the color codes are: red for the right ventricle, green for the myocardium, and blue for the left ventricle. From left to right, we show the results from TransUNet, Swin-UNet, AgileFormer, our SACNet with $L_{CD}$ (a combination of cross-entropy loss and Dice loss), our SACNet with $L_{CT}$ (our proposed dynamic loss), and the ground truth.
III-C Model Architectures
The overall architecture of our model is illustrated in Fig. 2. Like most medical image segmentation modeling methods [8][10], an encoder-decoder design is employed. Our framework contains four stages of adaptive spatial aggregation feature extraction encoders and a shared lightweight decoder. For an input RGB medical image with the shape of $H\times W\times 3$, we first use stem layers to reduce the spatial resolution. Subsequently, four feature maps ($X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$) in four different stages are generated and concatenated with the up-sampled features of the decoder through skip connections. The extracted features from the last stage ($X_{4}$) are also passed through a bottleneck layer, where the channel size is adjusted to gather more information while keeping the image size the same. The features from the bottleneck layer are propagated upward, up-sampled by a scale of two, and concatenated with the features from the encoder. We apply bilinear interpolation to generate the final segmentation maps, which pass through a softmax layer to classify each pixel.
III-D Training Loss
Recent 2D medical image segmentation methods combine the Dice loss with the cross-entropy loss as the final loss function. However, for multi-object segmentation tasks, the standard Dice loss, due to cosine similarity’s disregard for data distribution characteristics [25], treats each class equally, which tends to result in imbalanced segmentation. In this study, we tested most of the proposed loss functions for multi-class medical image segmentation and comprehensively considered the loss aggregation strategy. We propose a novel Continuity Dynamic Adjustment Loss Function (CTLoss) based on t-vMF Dice loss [25] and the cross-entropy loss function. This new loss function adaptively uses more compact similarities for easy classes and wider similarities for difficult classes, making it more effective in helping the model converge better and more efficiently for multi-object image segmentation.
The t-vMF Dice Loss is formulated as:
$\cos\theta_{i}=\sum_{i=1}^{n}A_{in}B_{in},$ (4)
@@ III-D Training Loss
The multi-class cross-entropy loss function is defined as follows:
$L_{CE}=-\sum_{i=1}^{N}\sum_{c=1}^{C}y_{ic}\log(\hat{y}_{ic}),$ (7)
where $N$ represents the total number of samples, and $C$ is the total number of classes. $y_{ic}$ denotes the true label for sample $i$ and class $c$ (if the sample belongs to class $c$, then $y_{ic}=1$, otherwise $y_{ic}=0$). $\hat{y}_{ic}$ is the predicted probability that sample $i$ belongs to class $c$.
$L_{CT}=\gamma L_{t-vMF}+(1-\gamma)L_{CE},$ (8)
where $\gamma$ is the weight coefficient, with $\gamma=0.6$ and $0.41-\gamma=$ 0.4 representing the weights for the adaptive t-vMF Dice loss and the cross-entropy loss, respectively.
IV Experiment
IV-A Dataset
Consistent with previous work, we validated the proposed method on two public medical image segmentation datasets: the Synapse multi-organ segmentation dataset and the Automated Cardiac Diagnosis Challenge (ACDC) dataset.
The Synapse dataset consists of 30 scans of eight abdominal organs from the MICCAI 2015 Multi-Atlas Abdominal Labeling Challenge. These eight organs include the left kidney, right kidney, aorta, spleen, gallbladder, liver, stomach, and pancreas. There are a total of 3779 axially enhanced abdominal clinical CT images. Following previous studies, we report the mean Dice Similarity Coefficient (DSC) and the 95% Hausdorff Distance (HD95) for the segmentation of abdominal organs using 18 training cases (2212 axial slices) and 12 validation cases.
The ACDC dataset consists of cine MR images acquired during breath-hold, covering the heart from the base to the apex of the left ventricle. Each scan was annotated with ground truth for the Left Ventricle (LV), Right Ventricle (RV), and Myocardium (Myo). Following previous studies, we report the mean DSC for 70 training cases (1930 axial slices), 10 validation cases, and 20 test cases.
IV-B Evaluation
We evaluate our model’s performance on the ACDC and Synapse multi-organ datasets using the DSC and the HD95 for the Synapse dataset. The DSC and HD95 are calculated according to the formulas as follows:
$DSC(X,Y)=\frac{2\cdot|X\cap Y|}{|X|+|Y|},$ (9)
$HD95(A,B)=\max\left\{\max_{a\in A}\min_{b\in B}\|a-b\|,\max_{b\in B}\min_{a\in A}\|b-a\|\right\},\small$ (10)
where $X$ represents the predicted segmentation mask, and $Y$ represents the ground truth segmentation mask. $a$ and $b$ are points in sets $A$ and $B$ respectively. $A$ represents the set of points in the predicted segmentation mask, and $B$ represents the set of points in the ground truth segmentation mask.
IV-C Setups
The model and experiments were implemented using PyTorch 2.2.0 and the CUDA toolkit 12.1, with all experiments conducted on an NVIDIA 3090 24G GPU. To improve computational efficiency, the resolution of all training, validation, and test images was resized to 224×224. We used a batch size of 18, an initial learning rate of $3\times 10^{-4}$, and trained the models for 400 epochs. All models were trained using a combination of t-vMF Dice Loss and cross-entropy losses, incorporating our proposed continuity dynamic adaptive algorithm. The AdamW optimizer was used, employing cosine learning rate decay. For better model initialization, we initialized the model parameters with weights pre-trained on ImageNet-1k. The DSC and the HD95 were used as metrics to evaluate segmentation accuracy. During training, the best validation weights were recorded to test the performance of our model.
TABLE II: Performance comparison with 2D methods in cardiac MRI segmentation using the ACDC dataset. The best result within each column is highlighted in bold, and the second-best is underlined. Models marked with an asterisk () were implemented by us.
[…]
R50+UNet [9] 2021 87.10 80.63 94.92 87.55%
TransUNet [9] 2021 86.67 87.27 95.18 89.71%
MISSFormer [14] 2021 89.55 88.04 94.99 90.86%
Swin-UNet [13] 2022 85.77 84.42 94.03 88.07%
MT-UNet [32] 2022 86.64 89.04 95.62 90.43%
TransCASCADE [33] 2023 89.14 90.25 95.50 91.63%
Parallel MERIT [34] 2024 90.87 90.00 96.08 92.32%
AgileFormer [20] 2024 90.93 90.29 95.99 92.41%
SACNet+$L_{CD}$(Ours) 2024 91.05 89.96 96.01 92.34%
SACNet+$L_{CT}$(Ours) 2024 91.38 90.11 96.10 92.53%
TABLE III: Comparison of Ablation Studies: Only a single parameter is altered while keeping the others constant to observe the impact of each parameter on performance. D represents the DCNv3 block, F represents the FFN block, and S represents the Layer Scale and DropPath block.
@@ IV-C Setups
V Result
V-A Performance on Synapse Dataset

### v2

@@ Abstract
Multi-organ segmentation in medical image analysis is crucial for diagnosis and treatment planning. However, many factors complicate the task, including variability in different target categories and interference from complex backgrounds. In this paper, we utilize the knowledge of Deformable Convolution V3 (DCNv3) and multi-object segmentation to optimize our Spatially Adaptive Convolution Network (SACNet) in three aspects: feature extraction, model architecture, and loss constraint, simultaneously enhancing the perception of different segmentation targets. Firstly, we propose the Adaptive Receptive Field Module (ARFM), which combines DCNv3 with a series of customized block-level and architecture-level designs similar to transformers. This module can capture the unique features of different organs by adaptively adjusting the receptive field according to various targets. Secondly, we utilize ARFM as building blocks to construct the encoder-decoder of SACNet and partially share parameters between the encoder and decoder, making the network wider rather than deeper. This design achieves a shared lightweight decoder and a more parameter-efficient and effective framework. Lastly, we propose a novel continuity dynamic adjustment loss function, based on t-vMF dice loss and cross-entropy loss, to better balance easy and complex classes in segmentation. Experiments on 3D slice datasets from Synapse demonstrate that SACNet delivers superior segmentation performance in multi-organ segmentation tasks compared to several existing methods.
@@ I Introduction
Multi-organ medical image segmentation, which simultaneously provides clearer visualization of anatomical and pathological structures across multiple organs, significantly enhances diagnostic efficiency and accuracy [1]-[2].
@@
However, it remains challenging due to several difficulties: (1) Variability across target categories. Not only do different organs within the same dataset vary significantly in size, shape, and texture, but even different slices of the same organ can exhibit considerable differences. This variability makes it challenging for segmentation models to generalize effectively, as they may overfit to specific features seen during training and struggle with unseen morphological variations. (2) Interference from complex backgrounds. The complexity of surrounding anatomical structures and the presence of varying textures, intensities, and overlapping tissues introduce substantial challenges. These intricate backgrounds can lead to ambiguous boundaries and obscure the precise contours of target organs, compromising the model’s ability to accurately segment organs and often resulting in incomplete or inaccurate segmentation outcomes.
To address these challenges in multi-organ medical image segmentation, recent studies [3]-[16] have employed various methods, including standard, dilated, and large-kernel convolutions, transformers, and hybrid architectures, to improve segmentation accuracy. Dilated and large-kernel convolutions expand the receptive field, while transformers [9] utilize self-attention to capture long-range dependencies. However, these models often lack adaptability to the specific structures of different organs, leading to potential segmentation inaccuracies. Some researchers [17]-[20] have incorporated deformable convolutions and deformable attention transformers to address these issues, though challenges such as sampling flexibility, computational complexity, and data requirements remain. Additionally, enhancements to the Dice loss function [21]-[26] have been proposed to handle segmentation imbalances, but limitations persist.
We propose the SACNet framework, which features an Adaptive Receptive Field Module (ARFM), a widenet strategy, and a dynamic continuity adjustment loss function:
  • ARFM: To handle varying target categories and complex backgrounds, we combine DCNv3 with grouped convolutions, ensuring each segmentation target has an adaptive receptive field. Additional elements like Feed-Forward, Layer Scale, and DropPath blocks enhance stability and geometric perception.
  • WideNet Strategy: To reduce computational load, we share DCNv3 block weights between encoder and decoder, allowing the network to be wider instead of deeper, achieving a lightweight and efficient design.
  • Continuity Dynamic Adjustment Loss Function (CTLoss): Combining t-vMF Dice loss and cross-entropy loss, CTLoss adjusts based on the Intersection Over Union (IOU) after each epoch, handling imbalanced data by dynamically adjusting the similarities for different classes.
Our method, SACNet, outperforms existing approaches in multi-organ segmentation tasks, as demonstrated by state-of-the-art results and thorough ablation studies.
Figure 1: Overview of the SACNet architecture.
II METHOD
II-A Encoder
Recent developments in large-scale CNN models [27] highlight the potential of Deformable Convolution v3 (DCNv3) for downstream visual tasks. In SACNet, the ARFM module leverages the pre-trained DCNv3 block as a core component, integrating it with Transformer-like designs to form the spatially adaptive pyramid feature encoder. The input image, sized 3×224×224, begins with a Stem Layer for initial processing. The encoder consists of four stages with embedding sizes of 112, 224, 448, and 896, each containing a series of ARFM modules and down-sampling layers. Each ARFM module includes five blocks: DCNv3 for adaptive receptive fields, Layer Norm, Feed-Forward block for pattern extraction, and Layer Scale and DropPath to stabilize training.
We emphasize the third encoder stage, using an ARFM module ratio of [4, 4, 21, 4] for optimal computation distribution. Skip connections are employed to maintain gradient flow. The encoder produces four feature maps of size $\{1\times(112\times 2^{i})\times\frac{H}{2^{2+i}}\times\frac{W}{2^{2+i}}\}$, where $i=\{0,1,2,3\}$, which are linked to the decoder via skip connections. The module sequence is illustrated in Fig.1.
II-B Decoder
The SACNet decoder is composed of four decoder blocks and one bottleneck block. Each block starts with layer normalization and upsamples the input by a factor of two to match the encoder outputs. The ARFM module refines features, followed by layer normalization, maintaining the same dimensions. The outputs are concatenated with the encoder outputs via skip connections. To handle the growing complexity due to DCNv3’s sampling points, we apply the WideNet [31] strategy, sharing DCNv3 projection weights between encoder and decoder to reduce parameter load. This allows SACNet to expand in width while limiting depth. After upsampling, images are resized to $H\times W\times N$ (where $N$ is the number of classes) for final segmentation, followed by Softmax for pixel classification. The bottleneck block functions similarly to other decoder blocks but focuses more on channel information without altering image size. No additional techniques like residual connections [29] or deep supervision [30] are used in the decoder.
@@ II-B Decoder
TABLE II: Comparison of Ablation Studies: Only a single parameter is altered while keeping the others constant to observe the impact of each parameter on performance. S represents the Layer Scale and DropPath block.
@@
Figure 2: Qualitative comparisons of the segmentation performance of our approach alongside other segmentation methods. The first two rows display Synapse dataset results with color codes: yellow (aorta), green (gallbladder), orange (left kidney), cyan (right kidney), plum (liver), purple (pancreas), indigo (spleen), and crimson (stomach).
II-C Training Loss
In this study, we tested most of the proposed loss functions for multi-class medical image segmentation and comprehensively considered the loss aggregation strategy. We propose a novel Continuity Dynamic Adjustment Loss Function (CTLoss) based on t-vMF Dice loss [25] and the cross-entropy loss function. This new loss function adaptively uses more compact similarities for easy classes and wider similarities for difficult classes, making it more effective in helping the model converge better and more efficiently for multi-object image segmentation.
The t-vMF Dice Loss is formulated as:
$\cos\theta_{i}=\sum_{i=1}^{n}A_{in}B_{in},$ (1)
$L_{CT}=\gamma L_{t-vMF}+(1-\gamma)L_{CE},$ (4)
where $L_{CE}$ is the cross-entropy loss function, $\gamma$ is the weight coefficient, with $\gamma=0.6$ and $0.41-\gamma=$ 0.4 representing the weights for the adaptive t-vMF Dice loss and the cross-entropy loss, respectively.
III Experiment
III-A Dataset and Evaluation
The Synapse dataset consists of 30 scans of eight abdominal organs from the MICCAI 2015 Multi-Atlas Abdominal Labeling Challenge. These eight organs include the left kidney, right kidney, aorta, spleen, gallbladder, liver, stomach, and pancreas. We report the mean Dice Similarity Coefficient (DSC) and the 95% Hausdorff Distance (HD95) for the segmentation of abdominal organs using 18 training cases (2212 axial slices) and 12 validation cases.
III-B Setups
The model and experiments were implemented using PyTorch 2.2.0 and the CUDA toolkit 12.1, with all experiments conducted on an NVIDIA 3090 24G GPU. To improve computational efficiency, the resolution of all training, validation, and test images was resized to 224×224. We used a batch size of 18, an initial learning rate of $3\times 10^{-4}$, and trained the models for 400 epochs.For better model initialization, we initialized the model parameters with weights pre-trained on ImageNet-1k.
IV Result
IV-A Performance on Synapse Dataset